OpenStudy (anonymous):

log (20x)= 3+ log (x-5) answer: (250)/(49) I have to assume the log base is 10 right? I've got this so far: log(20x/x-5)=10^3 But this does not seem right. Can anyone help? :O

OpenStudy (anonymous):

that last line should be: \(\large \frac{20x}{x-5}=10^3 \)

OpenStudy (anonymous):

Yes. That is what I have so far. What do I do from here? D:

OpenStudy (anonymous):

Solve for x. 20x/x-5 = 1000 20x = 1000x - 5000 5000 = 980x x = 5000/980 = 250/49

OpenStudy (anonymous):

Ok!! So I did it right! Phew! Thank you so so so so so much!!! :)

OpenStudy (anonymous):

np